4. With what initial velocity will a body moving along a vertical line have to be thrown if after 5 sec it is to be 50 ft above its starting place?


1
Expert's answer
2021-09-14T09:36:27-0400

Given:

"t=5\:\rm s"

"d=50\:\rm ft"


The vertical distance

"d=v_0t-gt^2\/2"

"v_0=\frac{2d+gt^2}{2t}\\
=\frac{2*50+32*5^2}{2*5}=90\:\rm ft/s=27.4\: m/s"